OpenStudy (anonymous):

Can someone explain how to do it or give me a formula for this? When Isabella pushes a box of mass 5.25 kilograms with a force of 15.7 newtons, it accelerates by a rate of 2.50 meters/second^2. What is the force due to friction in this scenario? A) -2.60 newtons B) 5.50 newtons C) -7.75 newtons D) 13.5 newtons

OpenStudy (anonymous):

With no friction, the acceleration should be: a=15.7/5.25=3m/sec^2. However the acceleration is 2.5, that means there is an opposite force (friction) whose value is (3-2.5)·5.25=2.6 Newtons. As it goies against the force Isabella is pushing, then negative sign

OpenStudy (***[isuru]***):

 Hi, With the use of Newton's second law; F = ma where m = mass a = acceleration F = force = (5.25*2.5) = 13.125 N So when 15.7 N is applied the box get affect by 13.125 N which means 1.575 N force has wasted during motion therefor F = 1.575 N (friction)